# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.1 Tangent bundle and vector field probabilities
$\mathcal{W}_{p}(e^{V}_{s}(\mu),e^{V}_{t}(\mu))\leq|t-s|\mathcal{M}_{1}(V[\mu]),$ (3.9)
for all $s,t\in[0,T]$. Let us now prove the second part of the statement. Suppose first $(M,g)=\mathbb{R}^{n}$ and let $(e_{i})_{i\in\llbracket 1,n\rrbracket}$ be the canonical base. Take $\mu\in\mathcal{W}_{p}(\mathbb{R}^{n})$, $x\in\mathbb{R}^{n}$, $i\in\llbracket 1,n\rrbracket$ and $h>0$. Then, there exists a bounded function $K:\mathcal{X}(M)\to\mathbb{R}$ such that
$\bar{V}[\mu](x+he_{i})$ $-\bar{V}[\mu](x)=\int_{\mathcal{X}(M)}\pi_{x+he_{i}}(X)-\pi_{x}(X)\operatorname{d}V[\mu](X)$ (3.10)
$=\int_{\mathcal{X}(M)}h\pi_{x}(\frac{\operatorname{d}X}{\operatorname{d}x_{i}})+K(X)(h^{2})\|X\|_{W^{2,\infty}(\mathbb{R}^{n})}\operatorname{d}V[\mu](X).$
@@
$\bar{V}[\mu](x+he_{i})-\bar{V}[\mu](x)-\int_{\mathcal{X}(M)}h\pi_{x}(\frac{\operatorname{d}X}{\operatorname{d}x_{i}})\operatorname{d}V[\mu](X)=_{h\to 0}O(h^{2}).$ (3.11)
Thus, $\frac{\operatorname{d}V[\mu]}{\operatorname{d}x_{i}}(x)=\int_{\mathcal{X}(M)}\pi_{x}(\frac{\operatorname{d}X}{\operatorname{d}x_{i}})\operatorname{d}V[\mu](X)$. Now, we readily show that $\overline{V}[\mu]\in W^{2,\infty}(M)$ and it follows that $\|\overline{V}[\mu]\|_{W^{2,\infty}(M)}\leq 3\mathcal{M}(V[\mu])$. Finally, an analogous proof in local coordinates yields the local result on a general Riemannian manifold $(M,g)$.
@@
$(\overline{V}[\mu]-\overline{V}[\nu])(x)$ $=\int_{\mathcal{X}(M)}\pi_{x}(X)\operatorname{d}(V[\mu]-V[\nu])(X)$ (3.12)
$\leq\int_{\mathcal{X}(M)}\|X\|_{\infty}\operatorname{d}(V[\mu]-V[\nu])(X)\leq\mathcal{W}_{1}(V[\mu],V[\nu])$
@@
$|\nabla\overline{V}[\mu]-\nabla\overline{V}[\nu]|(x)$ $\leq\int_{\mathcal{X}(M)}\|X\|_{W^{1,\infty}(M)}\operatorname{d}(V[\mu]-V[\nu])(X)$ (3.13)
@@
3.2 The parabolic Wasserstein–flow equation
@@
where $\square_{\mu(t)}^{V}\varphi=\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}_{X-\overline{V}[\mu(t)]}^{2}(\varphi)+\mathcal{L}_{X}(\varphi)\operatorname{d}V(\mu(t))$.
@@
With the notations from Definition 3.9, (3.14) reduces to
@@
$=\frac{1}{2}\int_{0}^{s}\int_{M}\int_{\mathcal{X}(M)}\mathcal{L}_{X-\overline{V}[\mu(t)]}^{2}(\varphi)\operatorname{d}V[\mu(t)]\operatorname{d}\mu(t)\operatorname{d}t$
$\qquad+\int_{0}^{s}\int_{M}\mathcal{L}_{\overline{V}[\mu(t)]}(\varphi)\operatorname{d}\mu(t)\operatorname{d}s.$
@@
$\int_{\mathcal{X}(M)}\mathcal{L}_{X}(\varphi)(x)\operatorname{d}V[\mu(t)](X)$ $=\int_{T_{x}M}\operatorname{d}_{x}\varphi(v)\operatorname{d}(\pi_{x})_{*}V[\mu(t)](v)$ (3.16)
$=\operatorname{d}_{x}\varphi(\int_{T_{x}M}v\operatorname{d}(\pi_{x})_{*}V[\mu(t)](v))$
@@
Note that a given curve $t\mapsto\mu(t)$ can be solution of (3.14) for several different vector field probabilities. In particular, one can without loss of generality assume that $V$ is symmetric.
Definition 3.11 (Symmetrisation of a vector field probability).
Let $V$ be a vector field probability on $\mathcal{W}_{p}(M)$. Let
$m_{Y}:\mathcal{X}(M)$ $\to\mathcal{X}(M)$ (3.17)
@@
for all $Y\in\mathcal{X}(M)$ and
$a:\mathcal{X}(M)$ $\to\mathcal{X}(M)$ (3.18)
@@
We define $SV$, the symmetrisation of $V$ as the vector probability
@@
Let $V$ be a vector field probability and $t\mapsto\mu(t)$ be a narrowly continuous curve of probability measures that solves (3.14) for $V$ with initial value $\mu_{0}$. Then, $\mu(t)$ solves (3.14) for $SV$.
3.3 Properties of solutions of the MDE
Before solving (3.14), let us first determine some basic properties of its solutions. To alleviate notations, we shall often write $\overline{X}[\mu]$ instead of $X-\overline{V}[\mu]$.
@@ 3.3 Properties of solutions of the MDE
Assume $(M,g)$ is of bounded geometry. Let $B\in\mathbb{R}^{*}_{+}$ and $V$ be a vector field probability on $\mathcal{W}_{p}(M)$ with pp–th moments uniformly bounded by $B$. Let $\mu:[0,T]\to\mathcal{W}_{1}(M)$ be a narrowly continuous curve of probability measures that solves (3.14) for $V$ with initial value $\mu_{0}$. Then, $t\mapsto\mu(t)$ is Hölder–continuous with exponent $\tfrac{1}{2}$ in $\mathcal{W}_{1}(M)$.
@@
$\biggl{|}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu(u)]}\varphi(x)\operatorname{d}V[\mu(u)](X)\operatorname{d}\mu(u)(x)\biggr{|}$
$\qquad\leq B^{\prime}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})$
@@
Hence, Corollary 2.21 yields that there exists a 11–Lipschitz $\varphi_{s,t}\in\mathcal{C}^{\infty}_{c}(M)$ and $K>0$, such that
@@
Let $B\in\mathbb{R}^{*}_{+}$ and $V$ be a vector field probability on $\mathcal{W}_{p}(M)$ with pp–th moments uniformly bounded by $B$. Let $\mu:[0,T]\to\mathcal{W}_{p}(M)$ be a narrowly continuous curve of probability measures that solves (3.14) for $V$ with initial value $\mu_{0}$. Then, $\mu(t)$ is uniformly continuous in $\mathcal{W}_{p}(M)$. Furthermore, let $\varepsilon>0$ and $\mathfrak{R}(\varepsilon)>0$ be the minimal value such that for all $t\in[0,T],R\geq\mathfrak{R}(\varepsilon)$,
@@
  • For $k=1$, the property follows directly from Proposition 3.13 if $(M,g)$ is of bounded geometry. Otherwise, let $\varepsilon>0$ and $s,t\in[0,T]$. Then, (3.14) yields that there exists a constant $C_{1}$ independent of $t$ and $s$, such that for all $\varphi\in\mathcal{C}^{\infty}_{c}(M)$,
@@
    $\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu(t)\leq\varepsilon\quad\text{and}\quad\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu(s)\leq\varepsilon.$
    Let $0\leq\varphi\in C^{\infty}_{c}(M)$ be such that $\varphi(x)=d(x_{0},x)$ in $B(x_{0},R)\setminus B(x_{0},1)$, and $\|\nabla\varphi\|_{\infty}\leq 1$, $\|\nabla^{2}\varphi\|_{\infty}\leq 1\leq 1$ and $d(x_{0},x)\leq\varphi(x)\leq 1$ for all $x\in B(x_{0},1)$. Such a function exists for all $R>0$. Then,
@@
    and the $(\mathcal{M}_{1}(\mu(t)))_{t}$ are uniformly bounded.
@@
    $\qquad\leq C\bigl{(}\mathcal{M}_{k}(\mu(s))+\mathcal{M}_{k-1}(\mu(s))\bigr{)}\leq C(K_{k}+K_{k-1}),$
@@
    The narrow continuity of $t\mapsto\mu(t)$ and continuity of the $t\mapsto\mathcal{M}_{k+1}(\mu(t))$ combined imply continuity in $\mathcal{W}_{k+1}(M)$. In particular, the $(\mathcal{M}_{k+1}(\mu(t)))_{t}$ are uniformly bounded.
@@
Finally, suppose $M$ is bounded geometry. Let $q<p$ and let $B>0$ such that for all $t\in[0,T]$, $\mathcal{M}_{p}(\mu(t))\leq B$. Then, by Lemma 2.11, for $R_{\varepsilon}=\bigl{(}B^{p}\frac{p}{p-q}\bigr{)}^{\frac{1}{p-q}}\varepsilon^{\frac{1}{q-p}}$, we have $\int_{M\setminus B(x_{0},R_{\varepsilon})}d(x_{0},x)^{q}\operatorname{d}\mu(t)\leq\varepsilon$. By Corollary 2.21, there exists $C>0$, such that
@@
Therefore, if $|t-s|\leq\varepsilon^{\frac{2(p-1)}{p-q}}$, $\mathcal{W}_{p}(\mu(t),\mu(s))\leq(2+C)\varepsilon$. Therefore, $\mu(t)$ is Hölder–continuous in $\mathcal{W}_{q}(M)$.
@@ 3.4 Motivation
$\nu$ $\mapsto\int_{\mathcal{X}(M)}(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}})_{*}\nu\operatorname{d}V[\mu](X)$
@@
$\nu$ $\mapsto\int_{\mathcal{X}(M)}(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}})_{*}\nu\operatorname{d}V[\nu](X)$
@@
$\int_{M\times\mathcal{X}(M)}f\circ\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}(x)\operatorname{d}V[\nu]\otimes\nu(X,x)$ (3.33)
$\qquad=\int_{\mathcal{X}(M)}\int_{M}f\circ\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}(x)\operatorname{d}\nu(x)\operatorname{d}V[\nu](X).$
Thus, $\int_{\mathcal{X}(M)}(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}})_{*}\nu\operatorname{d}V[\nu](X)$, defined by
$\int_{M}f(x)\operatorname{d}\bigl{(}\int_{\mathcal{X}(M)}(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}})_{*}\nu\operatorname{d}V[\nu](X)(x)\bigr{)}$ (3.34)
$\qquad=\int_{\mathcal{X}(M)}\int_{M}f\circ\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}(x)\operatorname{d}\nu(x)\operatorname{d}V[\nu](X),$
@@
For motivation purposes, we keep notations short here. In the further course of this work, we shall introduce general subdivisions of $[0,T]$ with maximal step length $\tau$, define a $\mu^{\tau}$ for each of these subdivisions. Naturally we would want all these approximations to converge to a common limit for $\tau\to 0$. We shall see that this is indeed the case with suitable assumptions on $V$ and $(M,g)$.
This implicit definition of a solution of (MDE) might be intuitively easy to grasp, but is hard to handle from a mathematical point of view. To obtain the explicit integral equation (3.14) from the idea outlined above, we argue as follows.
@@
$\mu(t+\operatorname{d}t)=\int_{\mathcal{X}(M)}(\rho^{X-\overline{X}}_{\sqrt{\operatorname{d}t}})_{*}(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)\operatorname{d}V(\mu(t))(X)$ (3.35)
@@
$\frac{\mu(t+\operatorname{d}t)-\mu(t)}{\operatorname{d}t}$ $=\int_{\mathcal{X}(M)}\frac{(\rho^{X-\overline{X}}_{\sqrt{\operatorname{d}t}})_{*}(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)-\mu(t)}{\operatorname{d}t}\operatorname{d}V(\mu(t))(X)$ (3.36)
$=\int_{\mathcal{X}(M)}\frac{(\rho^{X-\overline{X}}_{\sqrt{\operatorname{d}t}})_{*}(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)-(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)}{\operatorname{d}t}$
$\qquad+\frac{(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)-\mu(t)}{\operatorname{d}t}\operatorname{d}V(\mu(t))(X).$
@@
$=\int_{M}\varphi\operatorname{d}\left(\int_{\mathcal{X}(M)}\frac{(\rho^{X-\overline{X}}_{\sqrt{\operatorname{d}t}})_{*}(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)-(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)}{\operatorname{d}t}\operatorname{d}V(\mu(t))(X)\right)$
$\qquad\qquad+\int_{M}\varphi\operatorname{d}\biggl{(}\frac{(\rho^{\overline{X}}_{\operatorname{d}t})_{*}\mu(t)-\mu(t)}{\operatorname{d}t}\biggr{)}.$
@@
$\int_{M}\varphi\operatorname{d}\biggl{(}\frac{\mu(t+\operatorname{d}t)-\mu(t)}{\operatorname{d}t}\biggr{)}$ $=\int_{M}\frac{1}{\operatorname{d}t}\int_{\mathcal{X}(M)}\varphi\circ\rho^{X-\overline{X}}_{\sqrt{\operatorname{d}t}}-\varphi\operatorname{d}V(\mu(t))(X)\operatorname{d}\mu(t)$ (3.38)
$+\frac{\operatorname{d}}{\operatorname{d}t}\int_{M}\varphi\circ\rho^{\overline{X}}_{\operatorname{d}t}-\varphi\operatorname{d}\mu(t).$
With this equation it becomes clear why we have set the time step in the first integral to $\sqrt{\operatorname{d}t}$ instead of $t$: the expression $\int_{\mathcal{X}(M)}\varphi\circ\rho^{X-\overline{X}}_{\sqrt{\operatorname{d}t}}-\varphi\operatorname{d}V(\mu(t))(X)$ looks like the discrete version of some elliptic operator of second order. In fact, this is no surprise, since these operators usually appear when it comes to model diffusion phenomena. Of course, many operators of second order admit this particular discrete version. However, here the most logical choice is the one in Definition 3.9. Making the expression in the integral on the left-hand continuous is straightforward, the second integral on the right-hand side would rather become $\int_{M}\mathcal{L}_{\overline{X}_{V(\mu(t)}}(\varphi)\operatorname{d}V(\mu(t))$, see Lemma 3.10.

### v2

@@ 3.1 Tangent bundle and probability vector–fields
$\mathcal{W}_{p}\bigl((\rho^{\overline{V}[\mu]}_{s})_{*}\mu,(\rho^{\overline{V}[\mu]}_{t})_{*}\mu\bigr)\leq|t-s|\mathcal{M}_{1}(V[\mu]),$ (3.9)
for all $s,t\in[0,T]$. Let us now prove the second part of the statement. Suppose first $(M,g)=\mathbb{R}^{n}$ and let $(e_{i})_{i\in\llbracket 1,n\rrbracket}$ be the canonical base. Take $\mu\in\mathcal{W}_{p}(\mathbb{R}^{n})$, $x\in\mathbb{R}^{n}$, $i\in\llbracket 1,n\rrbracket$ and $h>0$. Then, there exists a bounded function $K:\mathcal{X}^{2,\infty}(M)\to\mathbb{R}$ such that
$\overline{V}[\mu](x+he_{i})$ $-\overline{V}[\mu](x)=\int_{\mathcal{X}^{2,\infty}(M)}\pi_{x+he_{i}}(X)-\pi_{x}(X)\operatorname{d}V[\mu](X)$ (3.10)
$=\int_{\mathcal{X}^{2,\infty}(M)}h\pi_{x}\biggl(\frac{\operatorname{d}X}{\operatorname{d}x_{i}}\biggr)+K(X)(h^{2})\|X\|_{W^{2,\infty}(\mathbb{R}^{n})}\operatorname{d}V[\mu](X).$
@@
$\overline{V}[\mu](x+he_{i})-\overline{V}[\mu](x)-\int_{\mathcal{X}^{2,\infty}(M)}h\pi_{x}\biggl(\frac{\operatorname{d}X}{\operatorname{d}x_{i}}\biggr)\operatorname{d}V[\mu](X)=_{h\to 0}O(h^{2}).$ (3.11)
Thus, $\frac{\operatorname{d}\overline{V}[\mu]}{\operatorname{d}x_{i}}(x)=\int_{\mathcal{X}^{2,\infty}(M)}\pi_{x}(\frac{\operatorname{d}X}{\operatorname{d}x_{i}})\operatorname{d}V[\mu](X)$. Now, we readily show that $\overline{V}[\mu]\in W^{2,\infty}(M)$ and it follows that $\|\overline{V}[\mu]\|_{W^{2,\infty}(M)}\leq 3\mathcal{M}(V[\mu])$. Finally, an analogous proof in local coordinates yields the local result on a general Riemannian manifold $(M,g)$.
@@
$(\overline{V}[\mu]-\overline{V}[\nu])(x)$ $=\int_{\mathcal{X}^{2,\infty}(M)}\pi_{x}(X)\operatorname{d}(V[\mu]-V[\nu])(X)$ (3.12)
$\leq\int_{\mathcal{X}^{2,\infty}(M)}\|X\|_{\infty}\operatorname{d}(V[\mu]-V[\nu])(X)\leq\mathcal{W}_{1}(V[\mu],V[\nu])$
@@
$|\nabla\overline{V}[\mu]-\nabla\overline{V}[\nu]|(x)$ $\leq\int_{\mathcal{X}^{2,\infty}(M)}\|X\|_{W^{1,\infty}(M)}\operatorname{d}(V[\mu]-V[\nu])(X)$ (3.13)
@@
3.2 The parabolic Wasserstein–flow equation (MDE)
@@
$\int_{M}\varphi\operatorname{d}\mu(s)-\int_{M}\varphi\operatorname{d}\mu(0)=\int_{0}^{s}\int_{M}\square_{\mu(t)}^{V}\varphi\operatorname{d}\mu(t)\operatorname{d}t,$ (MDEd)
where $\square_{\mu(t)}^{V}\varphi=\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}_{X-\overline{V}[\mu(t)]}^{2}(\varphi)+\mathcal{L}_{X}(\varphi)\operatorname{d}V[\mu(t)](X)$.
@@
With the notation from Definition 3.11, (MDEd) reduces to
@@
$=\frac{1}{2}\int_{0}^{s}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\mathcal{L}_{X-\overline{V}[\mu(t)]}^{2}(\varphi)\operatorname{d}V[\mu(t)]\operatorname{d}\mu(t)\operatorname{d}t$
$\hskip 18.49988pt+\int_{0}^{s}\int_{M}\mathcal{L}_{\overline{V}[\mu(t)]}(\varphi)\operatorname{d}\mu(t)\operatorname{d}s.$
@@
$\int_{\mathcal{X}^{2,\infty}(M)}\mathcal{L}_{X}(\varphi)(x)\operatorname{d}V[\mu(t)](X)$ $=\int_{T_{x}M}\operatorname{d}_{x}\varphi(v)\operatorname{d}(\pi_{x})_{*}V[\mu(t)](v)$ (3.15)
$=\operatorname{d}_{x}\varphi\biggl(\int_{T_{x}M}v\operatorname{d}(\pi_{x})_{*}V[\mu(t)](v)\biggr)$
@@
Note that a given curve $t\mapsto\mu(t)$ can be solution of (MDEd) for several different probability vector–fields. In particular, one can without loss of generality assume that $V$ is symmetric.
Definition 3.13 (Symmetrisation of a probability vector–field).
Let $V$ be a probability vector–field on $\mathcal{W}_{p}(M)$. Let
$m_{Y}:\mathcal{X}^{2,\infty}(M)$ $\to\mathcal{X}^{2,\infty}(M)$ (3.16)
@@
for all $Y\in\mathcal{X}^{2,\infty}(M)$ and
$a:\mathcal{X}^{2,\infty}(M)$ $\to\mathcal{X}^{2,\infty}(M)$ (3.17)
@@
We define $SV$, the symmetrisation of $V$ as the probability vector–field
@@
Let $V$ be a probability vector–field and $t\mapsto\mu(t)$ be a narrowly continuous curve of probability measures that solves (MDEd) for $V$ with initial value $\mu_{0}$. Then, $\mu(t)$ solves (MDEd) for $SV$.
3.3 Properties of solutions of (MDE)
Before solving (MDEd), let us first determine some basic properties of its solutions. To alleviate notation, we shall often write $\overline{X}[\mu]$ instead of $X-\overline{V}[\mu]$.
@@ 3.3 Properties of solutions of (MDE)
Assume $(M,g)$ is of bounded geometry. Let $B\in\mathbb{R}^{*}_{+}$ and $V$ be a probability vector–field on $\mathcal{W}_{p}(M)$ with pp–th moments uniformly bounded by $B$. Let $\mu:[0,T]\to\mathcal{W}_{1}(M)$ be a narrowly continuous curve of probability measures that solves (MDEd) for $V$ with initial value $\mu_{0}$. Then, $t\mapsto\mu(t)$ is Hölder–continuous with exponent $\tfrac{1}{2}$ in $\mathcal{W}_{1}(M)$.
@@
$\biggl|\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu(u)]}\varphi(x)\operatorname{d}V[\mu(u)](X)\operatorname{d}\mu(u)(x)\biggr|$
$\hskip 18.49988pt\leq B^{\prime}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})$
@@
Hence, Corollary 2.28 yields that there exist $K>0$ and a KK–Lipschitz $\varphi_{s,t}\in\mathcal{C}^{\infty}_{c}(M)$, such that
@@
Let $B\in\mathbb{R}^{*}_{+}$ and $V$ be a probability vector–field on $\mathcal{W}_{p}(M)$ with pp–th moments uniformly bounded by $B$. Let $\mu:[0,T]\to\mathcal{W}_{p}(M)$ be a narrowly continuous curve of probability measures that solves (MDEd) for $V$ with initial value $\mu_{0}$. Then, $\mu(t)$ is uniformly continuous in $\mathcal{W}_{p}(M)$. Furthermore, let $\varepsilon>0$ and $\mathfrak{R}(\varepsilon)>0$ be the minimal value such that for all $t\in[0,T],R\geq\mathfrak{R}(\varepsilon)$,
@@
  • For $k=1$, the property follows directly from Proposition 3.15 if $(M,g)$ is of bounded geometry. Otherwise, let $\varepsilon>0$ and $s,t\in[0,T]$. Then, (MDEd) yields that there exists a constant $C_{1}$ independent of $t$ and $s$, such that for all $\varphi\in\mathcal{C}^{\infty}_{c}(M)$,
@@
    $\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu(t)\leq\varepsilon\qquad\text{and}\qquad\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu(s)\leq\varepsilon.$
    Let $0\leq\varphi\in C^{\infty}_{c}(M)$ be such that $\varphi(x)=d(x_{0},x)$ in $B(x_{0},R)\setminus B(x_{0},1)$, and $\|\nabla\varphi\|_{\infty}\leq 1$, $\|\nabla^{2}\varphi\|_{\infty}\leq 1$ and $d(x_{0},x)\leq\varphi(x)\leq 1$ for all $x\in B(x_{0},1)$. Such a function exists for all $R>0$. Then,
@@
    and the $\bigl(\mathcal{M}_{1}(\mu(t))\bigr)_{t}$ are uniformly bounded.
@@
    $\hskip 18.49988pt\leq C\bigl(\mathcal{M}_{k}(\mu(u))+\mathcal{M}_{k-1}(\mu(u))\bigr)\leq C(K_{k}+K_{k-1}),$
@@
    The narrow continuity of $t\mapsto\mu(t)$ in combination with continuity of $t\mapsto\mathcal{M}_{k+1}(\mu(t))$ implies continuity in $\mathcal{W}_{k+1}(M)$. In particular, the $\bigl(\mathcal{M}_{k+1}(\mu(t))\bigr)_{t}$ are uniformly bounded.
@@
Finally, suppose $M$ is of bounded geometry. Let $q<p$ and let $B>0$ such that for all $t\in[0,T]$, $\mathcal{M}_{p}(\mu(t))\leq B^{p}$. Then, by Lemma 2.18, for $R_{\varepsilon}=\bigl(B^{p}\frac{p}{p-q}\bigr)^{\frac{1}{p-q}}\varepsilon^{\frac{1}{q-p}}$, we have $\int_{M\setminus B(x_{0},R_{\varepsilon})}d(x_{0},x)^{q}\operatorname{d}\mu(t)\leq\varepsilon$. By Corollary 2.28, there exists $C>0$, such that
@@
Therefore, if $|t-s|\leq\varepsilon^{\frac{2(p-1)}{p-q}}$, $\mathcal{W}_{q}(\mu(t),\mu(s))\leq(2+C)\varepsilon$. Therefore, $\mu(t)$ is Hölder–continuous in $\mathcal{W}_{q}(M)$.
@@ 3.4 Motivation
$\nu$ $\mapsto\int_{\mathcal{X}^{2,\infty}(M)}(\rho^{X-\overline{V}[\mu]}_{\sqrt{t}})_{*}\nu\operatorname{d}V[\mu](X)$
@@
$\nu$ $\mapsto\int_{\mathcal{X}^{2,\infty}(M)}(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}})_{*}\nu\operatorname{d}V[\nu](X)$
@@
$\int_{M\times\mathcal{X}^{2,\infty}(M)}f\circ\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}(x)\operatorname{d}V[\nu]\otimes\nu(X,x)$ (3.32)
$\hskip 18.49988pt=\int_{\mathcal{X}^{2,\infty}(M)}\int_{M}f\circ\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}(x)\operatorname{d}\nu(x)\operatorname{d}V[\nu](X).$
Thus, $\int_{\mathcal{X}^{2,\infty}(M)}\bigl(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}\bigr)_{*}\nu\operatorname{d}V[\nu](X)$, defined by
$\int_{M}f(x)\operatorname{d}\biggl(\int_{\mathcal{X}^{2,\infty}(M)}\bigl(\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}\bigr)_{*}\nu\operatorname{d}V[\nu](X)(x)\biggr)$ (3.33)
$\hskip 18.49988pt=\int_{\mathcal{X}^{2,\infty}(M)}\int_{M}f\circ\rho^{X-\overline{V}[\nu]}_{\sqrt{t}}(x)\operatorname{d}\nu(x)\operatorname{d}V[\nu](X),$
@@
For motivation purposes, we keep notation short here. In the further course of this work, we shall introduce general subdivisions of $[0,T]$ with maximal step length $\tau$, define a $\mu^{\tau}$ for each of these subdivisions. Naturally we would want all these approximations to converge to a common limit for $\tau\to 0$. We shall see that this is indeed the case with suitable assumptions on $V$ and $(M,g)$.
This implicit definition of a solution of (MDE) might be intuitively easy to grasp, but is hard to handle from a mathematical point of view. To obtain the explicit integral equation (MDEd) from the idea outlined above, we argue as follows.
@@
$\mu(t+\operatorname{d}t)=\int_{\mathcal{X}^{2,\infty}(M)}\bigl(\rho^{\overline{X}[\mu(t)]}_{\sqrt{\operatorname{d}t}}\bigr)_{*}\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)\operatorname{d}V[\mu(t)](X)$ (3.34)
@@
$\frac{\mu(t+\operatorname{d}t)-\mu(t)}{\operatorname{d}t}$ $=\int_{\mathcal{X}^{2,\infty}(M)}\frac{\bigl(\rho^{\overline{X}[\mu(t)]}_{\sqrt{\operatorname{d}t}}\bigr)_{*}\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)-\mu(t)}{\operatorname{d}t}\operatorname{d}V[\mu(t)](X)$ (3.35)
$=\int_{\mathcal{X}^{2,\infty}(M)}\frac{\bigl(\rho^{\overline{X}[\mu(t)]}_{\sqrt{\operatorname{d}t}}\bigr)_{*}\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)-\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)}{\operatorname{d}t}$
$\hskip 18.49988pt+\frac{\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)-\mu(t)}{\operatorname{d}t}\operatorname{d}V[\mu(t)](X).$
@@
$=\int_{M}\varphi\operatorname{d}\left(\int_{\mathcal{X}^{2,\infty}(M)}\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\bigl(\rho^{\overline{X}[\mu(t)]}_{\sqrt{\operatorname{d}t}}\bigr)_{*}\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)-\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)}{\operatorname{d}t}\operatorname{d}V[\mu(t)](X)\right)$
$\hskip 18.49988pt\hskip 18.49988pt+\int_{M}\varphi\operatorname{d}\biggl(\frac{\bigl(\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}\bigr)_{*}\mu(t)-\mu(t)}{\operatorname{d}t}\biggr).$
@@
$\int_{M}\varphi\operatorname{d}\biggl(\frac{\mu(t+\operatorname{d}t)-\mu(t)}{\operatorname{d}t}\biggr)$ $=\int_{M}\frac{1}{\operatorname{d}t}\int_{\mathcal{X}^{2,\infty}(M)}\varphi\circ\rho^{\overline{X}[\mu(t)]}_{\sqrt{\operatorname{d}t}}$ (3.37)
$\hskip 18.49988pt\hskip 18.49988pt-\varphi\operatorname{d}V[\mu(t)](X)\operatorname{d}\mu(t)$
$+\frac{\operatorname{d}}{\operatorname{d}t}\int_{M}\varphi\circ\rho^{\overline{V}[\mu(t)]}_{\operatorname{d}t}-\varphi\operatorname{d}\mu(t).$
With this equation it becomes clear why we have set the time step in the first integral to $\sqrt{\operatorname{d}t}$ instead of $t$: the expression $\int_{\mathcal{X}^{2,\infty}(M)}\varphi\circ\rho^{\overline{X}[\mu(t)]}_{\sqrt{\operatorname{d}t}}-\varphi\operatorname{d}V[\mu(t)](X)$ looks like the discrete version of some elliptic operator of second order. In fact, this is no surprise, since these operators usually appear when it comes to model diffusion phenomena. Of course, many operators of second order admit this particular discrete version. However, here the most logical choice is the one in Definition 3.11. Making the expression in the integral on the left-hand side continuous is straightforward, the second integral on the right-hand side would rather become $\int_{M}\mathcal{L}_{\overline{V}[\mu(t)]}(\varphi)\operatorname{d}\mu(t)$, see Lemma 3.12.